OpenStudy (anonymous):

Which is the 23rd term of the arithmetic series 5, 16, 27,

OpenStudy (amistre64):

what is the common difference between each term?

OpenStudy (anonymous):

6

OpenStudy (amistre64):

well, thats at least testable :) 5+6 = 16?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

i find it ironic that your doing an arithmetic series but havent mastered adding yet ...

OpenStudy (amistre64):

5+6 does not equal 16 5 + 11 = 16

OpenStudy (amistre64):

our common different is: 11 , we will need this to determine any term in the series

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

by adding any multiple of 11 to our first term, we can get to any other term in the series. an = 5 + 11(n-1)

OpenStudy (amistre64):

a32 is going to be our 32 term; so when n=32 ... a32 = 5 + 11(32-1) is our answer

OpenStudy (amistre64):

but then you want 23, not 32 :)

OpenStudy (amistre64):

a23 = 5 + 11(23-1) would suffice

OpenStudy (anonymous):

kay thanks

OpenStudy (amistre64):

good luck :)